# FlashDMoE: Fast Distributed MoE in a Single Kernel

### v2

@@ 21. Answer: [Yes]
  22. Justification: All reported results in the evaluation section were obtained as the average of 150 executions.
@@ 50. Answer: [N/A]
Appendix A Related Work
Computation-Communication Overlap and Kernel Fusion. To reduce the communication overheads of synchronization in distributed DNN training, many research efforts have been focused on increasing the overlap of computation and communication. For generic Transformer-based models without MoE layers, many works [41, 42, 43, 44, 45, 46, 47, 48, 49] have provided insights and techniques to partition and schedule computation and communication operations, aimed at finer-grained overlapping. To address the challenges posed by AllToAll communication and expert parallelism in MoE training, Tutel [50] and FasterMoE [13] overlap AllToAll with expert computation. Lancet [51] additionally enables both non-MoE computation in forward pass and weight gradient computation in backward pass to be overlapped with AllToAll. Despite overlapping, the performance of these approaches is limited in practice due to blocking synchronous collective communication with barriers. In contrast, FlashDMoE fundamentally eliminates these inefficiencies with asynchronous, device-initiated data transfers overlapped with tiled computation all within a single kernel. FlashDMoE further differentiates itself from SOTA works like COMET [11] and DeepEP [1], which also use this form of kernel-initiated communication but at a coarse-grained granularity and without complete kernel fusion.
Appendix B Motivation Plots
@@
Figure 15: Straggler effect of synchronous AllToAll. $M\times N$ A100 or V100 denotes $N$ GPUs within a node across $M$ nodes. Every GPU communicates with every other GPU per AllToAll step. We capture the distribution of delay induced by stragglers across many steps. Actual Time $t_{a}$ denotes the fastest kernel execution time across all GPUs, conversely Total Time $t$ is the maximum recorded step time, while $Delay$ is the maximum difference between $t$ and $t_{a}$. Note $Delay$ is idle time.
@@
For any source process $p_{s}$, a valid index coordinate $i=(p*,r,b,e,c)$ satisfies the following:
@@
$i_{1}=(p_{1},r_{1},b_{1},e_{1},c_{1}),\>i_{1}=(p_{2},r_{2},b_{2},e_{2},c_{2})$
Note that for the message staging state, even though $i_{1}=i_{2}$ the resultant memory segments reside in different physical buffers resident in $p_{s_{1}}$ and $p_{s_{2}}$ respectively. Therefore, for this state, there are no conflicts as intra-process writes always have distinct $c_{j}$ coordinates, where $j\in\{0,C-1\}$. For inter-process transfers, we have two cases.
@@
Appendix D Task Implementation
@@
Appendix E Actors
E.1 Processor
1 begin
@@ 14             $lt\leftarrow 0$
17                   Aggregate locally observed task counts into $\text{Aggregate locally observed task counts into }lt$
@@ 23             $\mathbf{WarpInclusiveSum}(lt,qS,tasktally)$
25                   Repopulate $\text{Repopulate }rQ\text{ with ready processor ids}$
@@
27                         Starting at $\text{Starting at }rQ[qS]\text{, signal processors about task indices from }tqState$
@@ 17                         Enforce memory consistency before consuming packet
18                         Decode packet into a set of $\text{Decode packet into a set of }GEMM_{0}\text{ task descriptors using }X$
19                         Write task descriptors to $\text{Write task descriptors to }tQ$
@@ 33                         Enforce memory consistency before consuming packet
34                         Decode packet into a set of $\text{Decode packet into a set of }combine\text{ task descriptors using }T_{\phi},\>G_{\phi},O$
35                         Write task descriptors to $\text{Write task descriptors to }tQ$
@@ 36                         Notify Scheduler of decoded tasks
Appendix F Multi-Node Evaluation
F.1 Setup
In this experiment, we seek to evaluate FlashDMoE in the multi-node setting. We use 4 nodes, where each node comprises 4 A100 GPUs fully interconnected via NVLink. Across nodes, each GPU uses a single NIC providing 25 GB/s of bandwidth. We set the number of experts to be $16$ and assign each GPU to host only one, so the number of local experts is $1$. Note that we define MIV formally as follows:
$MIV=\frac{Tokens}{Experts}*local\_{experts}*precision*hidden\_size*2*n_{rg}$
where $n_{rg}$ is the number of remote peers and the multiplicative factor of $2$ accounts for communication rounds (dispatch and combine). $n_{rg}=12$ for this experiment.
F.2 Results
Figure 17: Multi-node Latency evaluation. Embbeding dimension is $1024$ and FFN intermediate size is $4096$. We define Maximal Incast Volume (MIV) as the worst case upper bound for data volume that a NIC receives in a single incast occurence.
We observe a sublinear increase in latency as we scale the number of tokens. However, we observe at $Tokens>2048$, that the application fails to terminate due to failure to receive expectant messages. We hypothesize this failure to be due to buffer overflow at the networking hardware layer as is common for applications that generate many and large messages [18] like our system. We note that this failure is addressable by tuning hardware configurations [52] but we consider this exploration as an exercise orthogonal to this work.
@@ Appendix G Implementation
Table 4: Implementation metrics of FlashDMoE.
@@ Appendix H FP16 Memory Throughput
Figure 18: Here, we report the total A100 memory throughput for both FP16 (top) and FP32 (bottom) variants of FlashDMoE. Notably, the FP16 implementation issues approximately 2×2$\times$2 × more shared memory instructions compared to its FP32 counterpart under identical workloads. We attribute this inefficiency to suboptimal shared memory layouts in FlashDMoE when operating on half-precision data. While this bottleneck is addressable through improved layout strategies, we leave its resolution to future work due to time constraints.

### v3

@@ 21. Answer: [Yes]
  22. Justification: All reported results in the evaluation section were obtained as the average of 32 executions preceded by 32 warmup runs.
@@ 50. Answer: [N/A]
Appendix A Supplementary Motivation
Figure 13: Overlapped Schedule (bottom) showing how idle time from the sequential schedule (top) is repurposed for computation. FlashMoE implements the overlapped schedule.
@@ Appendix A Supplementary Motivation
Figure 14: Straggler effect of synchronous AlltoAll. $M\times N$ A100 or V100 denotes NN GPUs within a node across $M$ nodes. Every GPU communicates with every other GPU per AlltoAll step. We capture the distribution of delay induced by stragglers across many steps. Actual Time $t_{a}$ denotes the fastest kernel execution time across all GPUs, conversely Total Time $t$ is the maximum recorded step time, while $Delay$ is the maximum difference between $t$ and $t_{a}$. Note DelayDelay is idle time.
In Figure 14, we present empirical cumulative and raw distributions of AlltoAll kernel runtime from distributed training of a 1.3B GPT-3 MoE model across 32 A100 and 8 V100 GPUs. We use this result to motivate the severity and prevalence of straggler effects. In Figure 14(b), we observe P95 communication performance degradation of 1.32X when compared to the mean actual kernel time. This performance reduction is rather tame as the underlying hardware is a supercomputer well-tuned against “software jitter” [nerscNetworkNERSC]. However, we observe a more severe p95 performance loss of 11X in a single-node Virtual Machine (VM). In line with prior HPC works [1639320, 10.1145/3545008.3545056], we argue that obviating the inherent barrier in this synchronous collective communication would allow GPUs to repurpose this observed idle time for downstream computation as depicted in Figure 13.
Table 2: Straggler Delay within Synchronous All-to-All communication. We capture the distribution of delay induced by stragglers across many steps. Let Actual Time $t_{a}$ denote the fastest kernel execution time across all GPUs, and Total Time $t$ be the maximum recorded step time. We define DelayDelay as the maximum difference between $t$ and $t_{a}$. Note DelayDelay is idle time. For the 1x8 V100, we profile 1750 steps and 600 steps for the 8x4 A100. See Figure 14 for the raw distribution.
System # Nodes # GPUs Median p95
Commercial VM (V100) 1 8 3.1x 11.4x
Supercomputer (A100) 8 32 1.09x 1.32x
Appendix B Related Work
Computation-Communication Overlap and Kernel Fusion. To reduce the communication overheads of synchronization in distributed DNN training, many research efforts have been focused on increasing the overlap of computation and communication. For generic Transformer-based models without MoE layers, many works [coconet, decomposition, centauri, t3, megascale, co2, syndicate, ccfuser, fused] have provided insights and techniques to partition and schedule computation and communication operations, aimed at finer-grained overlapping. To address the challenges posed by AlltoAll communication and expert parallelism in MoE training, Tutel [tutel] and FasterMoE [fastermoe] overlap AlltoAll with expert computation. Lancet [lancet] additionally enables both non-MoE computation in forward pass and weight gradient computation in backward pass to be overlapped with AlltoAll. Despite overlapping, the performance of these approaches is limited in practice due to blocking synchronous collective communication with barriers. In contrast, FlashMoE fundamentally eliminates these inefficiencies with asynchronous, device-initiated data transfers overlapped with tiled computation all within a single kernel. FlashMoE further differentiates itself from SOTA works like COMET [comet] and DeepEP [deepep], which also use this form of kernel-initiated communication but at a coarse-grained granularity and without complete kernel fusion.
@@ Appendix B Related Work
For any source process $p_{s}$, a valid index coordinate i=(p,r,b,e,c)i=(p*,r,b,e,c) satisfies the following:
@@
$i_{1}=(p_{1},r_{1},b_{1},e_{1},c_{1}),\>i_{2}=(p_{2},r_{2},b_{2},e_{2},c_{2})$
Note that intra-process writes always have distinct $c_{j}$ coordinates, where $j\in\{0,C-1\}$. For inter-process transfers, we have two cases.
@@
Appendix D Memory Overhead
We measure the GPU memory required for the symmetric tensor $L$ and runtime bookkeeping state of FlashMoE. The memory overhead primarily depends on the tile size, expert capacity ($EC$), and the number of experts ($E$). Table 3 summarizes the memory overhead across recent MoE models [moonlight, grok, snowflake-arctic, qwen3, mixtral, deepseek] during inference, showing that FlashMoE maintains a modest and predictable memory footprint. In particular, the symmetric tensor ($ST$) accounts for at most 2.15% additional memory relative to the total inference memory requirements.
Table 3: Memory overhead of FlashMoE (tile size $bM=128$, $Size(T)=\text{Tokens}\times 4\text{KB}$).
Model Params S E H I ST (GB) Model (GB) Overhead (%)
Moonlight-16B-A3B 16B 8K 64 2K 1.38K 0.25 59 0.49
Grok-1 314B 8K 8 6K 32K 0.75 1169 0.15
Snowflake-Arctic 479B 4K 128 7K 4.75K 1.75 1784 0.12
Qwen3-235B-A22B 235B 40K 128 4K 1.5K 3.00 875 0.38
Mixtral 8x7B 47B 32K 8 4K 14K 2.00 175 2.15
DeepSeek-V3 685B 160K 256 7K 2K 1.50 2551 0.11
Appendix E Task Implementation
@@
Appendix F Actors
F.1 Processor
1begin
@@ 14    $lt\leftarrow 0$
17       $\text{Aggregate locally observed task counts into }lt$
@@ 23    $\mathbf{WarpInclusiveSum}(lt,qS,tasktally)$
25       $\text{Repopulate }rQ\text{ with ready processor ids}$
@@
[…]
@@ 17          Enforce memory consistency before consuming packet
[…]
19          $\text{Write task descriptors to }tQ$
@@ 33          Enforce memory consistency before consuming packet
[…]
35          $\text{Write task descriptors to }tQ$
@@ Appendix G Implementation
Table 4: Implementation metrics of FlashMoE.
@@ Appendix H FP16 Memory Throughput
Figure 16: Here, we report the total A100 memory throughput for both FP16 (top) and FP32 (bottom) variants of FlashMoE. Notably, the FP16 implementation issues approximately $2\times$ more shared memory instructions compared to its FP32 counterpart under identical workloads. We attribute this inefficiency to suboptimal shared memory layouts in FlashMoE when operating on half-precision data. While this bottleneck is addressable through improved layout strategies, we leave its resolution to future work.